### v1

@@ Abstract
With significant advancements in Transformers and large language models (LLMs), natural language processing (NLP) has extended its reach into many research fields due to its enhanced capabilities in text generation and user interaction. One field benefiting greatly from these advancements is cybersecurity. In cybersecurity, many parameters that need to be protected and exchanged between senders and receivers are in the form of text and tabular data, making NLP a valuable tool in enhancing the security measures of communication protocols. This survey paper provides a comprehensive analysis of the utilization of Transformers and LLMs in cyber-threat detection systems. The methodology of paper selection and bibliometric analysis is outlined to establish a rigorous framework for evaluating existing research. The fundamentals of Transformers are discussed, including background information on various cyber-attacks and datasets commonly used in this field. The survey explores the application of Transformers in intrusion detection systems (IDSs), focusing on different architectures such as Attention-based models, LLMs like bidirectional encoder representations from Transformers (BERT) and generative pre-trained Transformer (GPT), CNN/LSTM-Transformer hybrids, emerging approaches like Vision Transformers (ViTs), among others. Furthermore, it explores the diverse environments and applications where Transformers and LLMs-based IDS have been implemented, including computer networks, internet-of-things (IoT) devices, critical infrastructure protection, cloud computing, software-defined networking (SDN), as well as in autonomous vehicles (AVs). The paper also addresses research challenges and future directions in this area, identifying key issues such as interpretability, scalability, and adaptability to evolving threats, and more. Finally, the conclusion summarizes the findings and highlights the significance of Transformers and LLMs in enhancing cyber-threat detection capabilities, while also outlining potential avenues for further research and development.
@@ 1 Introduction
Current approaches detect pattern changes induced by attacks by learning the normal pattern of a sequence, typically extracted solely from network data frames like log files and host IDs. Consequently, if the target sequence for detection contains only a minimal number of attacks, the deviation from the normal pattern will be subtle, posing challenges for detection. Therefore, there is a need for a new detection method to address such scenarios. While recurrent neural network (RNN) and long short-term memory (LSTM) models are commonly employed to grasp time series characteristics [9, 10], they suffer from decreased performance as sequences lengthen. Given the inherently sequential nature of network traffic data, new sequential AI models, like the Transformer network model [11, 12], present a logical solution to this issue. Transformers and their variants, leveraging self-Attention mechanisms [11], have excelled in natural language processing (NLP) tasks, including text classification, dialogue recognition, and machine translation. Inspired by the Transformer’s prowess in handling ordered data sequences, researchers have explored its applicability in intrusion and anomaly detection, demonstrating its effectiveness in various scenarios. Since network intrusion typically unfolds over time, most existing models lack the capability to capture time series features, resulting in information loss. The Attention mechanism of the Transformer, however, can effectively learn the temporal correlation of network intrusion data, thereby enhancing the accuracy of network intrusion detection.
The potential application of these advancements lies in their utilization with various cyber-data, including computer logs and network packets, can be structured as text and conceptualized as distinct languages. A modest yet expanding body of literature has showcased successful applications of NLP to cyber data, such as identifying red team attack within authentication logs. Once attacks are accurately identified, straightforward responses like quarantining or shutting down affected systems become feasible. Furthermore, leveraging windows event logs enables each computer to autonomously execute its response without necessitating feedback from other network sources. This autonomy is especially valuable in tactical networks, where connections may be unreliable and exhibit low throughput [13]. Transformers are main component of large language model (LLM) models due to their ability to capture complex dependencies in sequential data. For example, generative pre-trained transformer (GPT) [14] and bidirectional encoder representations from transformers (BERT) models [12], use Transformers to generate coherent and contextually relevant text. By leveraging Attention mechanisms, Transformers enable LLM systems to understand and manipulate data at a more nuanced level, leading to more realistic and diverse outputs in tasks such as text generation, image generation, and speech synthesis [15]. Thus, the LLM could be adapted at learning the benign patterns found within network traffic sequences, making it well-suited to detect alterations in network traffic patterns induced by even a minimal number of attacks [16].
@@
This survey aims to summarize the cutting-edge IDS based on Transformers and LLM advances in AI techniques, which offer promising alternatives to address the weaknesses of existing ML and DL-based IDS state-of-the-art (SOTA) strategies. By reading this survey, the author aims to assist researchers in understanding the concept of enhancing IDS with Transformers and LLMs, as well as shedding light on future directions and perspectives.
@@ 1.1 Existing reviews and our contributions
  • The survey offers an extensive taxonomy of different Transformer models applicable to strengthen various IDS techniques, including Attention, CNN/LSTM-Transformer, vision Transformer (ViT), GAN-Transformer, GPT, BERT and their variants, among others.
  • It presents a detailed classification of various Attention mechanisms and their possible configurations with traditional DL models, such as convolution neural network (CNN).
@@ 1.3 Survey methodology
To identify and review existing studies on Transformers and LLM-based ASR, a comprehensive search was conducted across several leading publication databases renowned for their high-quality scientific research. The primary search was carried out in Scopus, which systematically includes databases such as Web of Science, Elsevier, IEEE, ACM Digital Library, Wiley, and IET Digital Library, among others.
Articles published between 2017 and 2024 were given priority. However, older publications were also considered when necessary to provide a historical context, dataset, metrics, etc. The survey focused on computer science and engineering studies from databases including IEEE Xplore, ScienceDirect, Wiley, Springer, and Taylor & Francis. Additionally, considering the recent and trending nature of the topic, high-quality pre-prints from arXiv, SSRN, and TechRxiv were selected. Only articles written in English were included in the final analysis. Key search terms used in the abstract, article title, and keywords led to the formulation of the following query:
@@
The symbols & and || signify the logical operations AND and OR, respectively. Due to the overlap of the term "Transformers" with electrical network equipment, a manual filtering process was applied to exclude irrelevant papers. The final number of selected papers on IDS based purely on Transformers and LLMs is 102. Figure 1 provides more details about the involved research papers in terms of paper type, domain conducted, and distribution of papers based purely on Transformers and LLMs. It is obvious that Transformers and LLM-based IDS has emerged as a recent and rapidly growing field, with a notable surge in publications starting from 2022. Moreover, the inclusion of 27 pre-prints among the 148 cited papers underscores the trending nature and active research pursuit in this domain.
Figure 1: Bibliographic Statistics: (a) Analysis of research field and article type distribution from all references (156 papers). (b) Statistical distribution of 102 research papers on Transformers and LLMs for IDS (2017-2024).
@@ 2.2 Attack categories, datasets, and metrics
In-Vehicle 4 8.69 M 1:11 Flooding, fuzzy, Spoofing, Replay Yes AV IDS [37] Yes [^2] https://ieee-dataport.org/open-access/car-hacking-attack-defense-challenge-2020-dataset
@@ 3 Taxonomy of Transformers in IDS
CNN-Transformer Used for combining spatial feature extraction with sequential modeling Captures both local and global patterns in data, improving detection accuracy Computationally intensive due to combining two complex models
BERT Applied for analyzing network traffic data for anomaly detection Leverages bidirectional context for better understanding of data patterns Requires large amounts of data for fine-tuning and significant computational resources
RoBERTa Fine-tuned for anomaly detection in network traffic using robustly optimized BERT approach Improved performance over BERT due to enhanced training techniques Even higher computational requirements and data needs compared to BERT
GPT Generates responses based on network behavior patterns for identifying anomalies Effective in generating realistic sequences, useful for simulation and detection May struggle with understanding context in bidirectional sense compared to BERT
Vision Transformer Adapts vision-based Transformer models for network traffic visualization and anomaly detection Excellent at capturing intricate patterns in high-dimensional data High computational cost and requires large datasets for training
Self-Attention Focuses on relevant parts of the input data for anomaly detection Improves detection by focusing on the most important parts of the data Can be computationally expensive and may require extensive tuning
Multi-head Attention Enhances self-Attention by using multiple Attention heads for better anomaly detection Provides multiple perspectives on the data, improving detection robustness Increases model complexity and computational requirements
GAN-Transformer Combines generative adversarial networks with Transformers for detecting and simulating anomalies Generates realistic anomalies, enhancing detection robustness High complexity and requires extensive training time
3.1 Attention-based methods
@@
In this mechanism, $Q$, $K$, and $V$ represent queries, keys, and values, respectively. Queries are used to direct the Attention mechanism’s focus. Keys are part of the input data used to calculate Attention weights, indicating the relevance of different parts of the input data. Values are the actual content that, after the application of Attention weights, are aggregated to form the output of the Attention mechanism. Additionally, $d_{k}$ denotes the dimension of the keys. This approach enables the model to concentrate on the most relevant parts of the data, thereby enhancing the precision and efficiency of security systems through more accurate threat identification and reduced false positives. Figure 4 summarizes the existing Attention categories, and their benefits.
Figure 4: Visual guide to attention types in seq2seq models, illustrating how focus shifts with input and output, along with their respective related works. The primary components of self-attention and its respective multi-head attention (MHA) are scaled dot-product and global (soft) attentions. Hierarchical attention constructs multiple layers of attention. Local hard attention is the main component of Sparse Transformer and Linformer. Cross-attention is a key component of encoder-decoder models and ViT Transformers. Conditional attention is central to many LLM models such as BERT, GPT variants, and T5. CANINE attention is unique to the CANINE Transformer, where tokenization principles are not adhered to or realized.
@@ 3.2 CNN/LSTM-Transformer-based methods
A CNN-Transformer combines the strengths of CNNs and Transformers, two powerful neural network architectures used in DL. CNNs excel in processing structured grid data like images, efficiently capturing local dependencies through convolutional operations. They are renowned for their ability to detect features and patterns at various levels of abstraction, making them pivotal in computer vision (CV) tasks. On the other hand, Transformers leverage self-Attention mechanisms to model long-range dependencies, excelling in handling sequential data such as text and time series. They have revolutionized NLP and are increasingly applied across various domains. By integrating CNNs for effective feature extraction and spatial hierarchy with Transformers’ capability to model complex data relationships, CNN-Transformers aim to harness the advantages of both architectures. This synergy enables more robust and efficient learning, particularly in tasks requiring Attention to both local and global data structures, such as advanced image recognition, object detection, and beyond. The possible configurations of CNN-Transformers are illustrated in Figure 6. CNN-Transformers, an innovative architectural fusion, have recently garnered Attention in the realm of IDS. This advanced merger harnesses the strengths of CNNs and Transformers to enhance detection accuracy and efficiency. For example, in [64], feature subsets covering multiple spaces are developed using CNN to enhance the spatial distribution of samples. Subsequently, the Transformer is employed to establish connections between features and to determine essential attributes, including the temporal and detailed aspects of the features, culminating in the successful detection of intrusion activities. Similarly, the work in [65] presents a hybrid neural network model aiming to improve feature extraction and detection effectiveness in traffic analysis. It combines dense CNNs and Transformers for feature fusion and time sequence extraction. Results on the CIC-IDS2018 dataset show 98% accuracy, surpassing existing models in performance metrics. Moving on, the study referenced in [66] employs 1D CNN and time series Transformer architectures to address the challenges posed by RNNs related to computational complexity and detection performance, particularly due to issues of information loss. The performance assessment on the CICDDoS2019 and CSE-CIC-IDS2018 datasets showed outstanding results, with the evaluation metrics mostly falling between 98.07% and 99.99%.
@@
Figure 6: Various configurations of CNN-Transformer mechanisms include: (a) CNN and Transformer alternated; (b) One-by-one CNN-Transformer feature fusion; (c) CNN combined with multi-scale Transformer feature fusion; (d) A sequence of CNN followed by a single Transformer; (e) A sequence of CNN followed by the last CNN and a single Transformer feature fusion; (f) Sequential integration of multi-scale CNN and Transformer feature fusion; (g) Specific-scale CNN with parallel Transformer feature fusion.
@@
Researchers have proposed numerous SOTA methods that combine CNN with LSTM to form hybrid schemes like CNN-LSTM, applied in various research fields including IDS [67], among others. These combinations have proven effective in enhancing overall accuracy. However, to handle long dependencies, researchers have increasingly either added Transformers or replaced LSTMs with Transformers. Wang et al. [39] suggest a model for detecting intrusions by integrating the pre-trained ResNet model, MHA-based Transformer, and bidirectional LSTM models into a single framework that captures both spatial and temporal characteristics of network traffic, leveraging their robust capabilities in learning data representations. This approach significantly reduces the time required to process complex, high-dimensional intrusion data through DL techniques. The objective is to address challenges in IDSs, such as inadequate feature extraction and imprecise classification with complex, nonlinear, and high-dimensional data. To ensure a balanced dataset, the synthetic minority oversampling technique (SMOTE) has been utilized for preprocessing.

### v2

@@ Abstract
With significant advancements in Transformers and large language models (LLMs), natural language processing (NLP) has extended its reach into many research fields due to its enhanced capabilities in text generation and user interaction. One field benefiting greatly from these advancements is cybersecurity. In cybersecurity, many parameters that need to be protected and exchanged between senders and receivers are in the form of text and tabular data, making NLP a valuable tool in enhancing the security measures of communication protocols. This survey paper provides a comprehensive analysis of the utilization of Transformers and LLMs in cyber-threat detection systems. The methodology of paper selection and bibliometric analysis is outlined to establish a rigorous framework for evaluating existing research. The fundamentals of Transformers are discussed, including background information on various cyber-attacks and datasets commonly used in this field. The survey explores the application of Transformers in intrusion detection systems (IDSs), focusing on different architectures such as Attention-based models, LLMs like BERT and GPT, CNN/LSTM-Transformer hybrids, and emerging approaches like Vision Transformers (ViTs), and more. Furthermore, it explores the diverse environments and applications where Transformers and LLMs-based IDS have been implemented, including computer networks, internet-of-things (IoT) devices, critical infrastructure protection, cloud computing, software-defined networking (SDN), as well as in autonomous vehicles (AVs). The paper also addresses research challenges and future directions in this area, identifying key issues such as interpretability, scalability, and adaptability to evolving threats, and more. Finally, the conclusion summarizes the findings and highlights the significance of Transformers and LLMs in enhancing cyber-threat detection capabilities, while also outlining potential avenues for further research and development.
@@ 1 Introduction
Current approaches detect pattern changes induced by attacks by learning the normal pattern of a sequence, typically extracted solely from network data frames like log files and host IDSs. Consequently, if the target sequence for detection contains only a minimal number of attacks, the deviation from the normal pattern will be subtle, posing challenges for detection. Therefore, there is a need for a new detection method to address such scenarios. recurrent neural network (RNN) and long short-term memory (LSTM) models are commonly employed to grasp time series characteristics [9, 10], it is pivotal to recognize that these models excel in capturing time dependencies within sequential data, a critical feature for log file analysis. Unlike traditional static data analysis techniques, RNNs and LSTMs can process continuous streams of data, adjusting their internal state based on the latest inputs. This dynamic adjustment is essential for detecting anomalies in log files where patterns evolve and new attack vectors emerge over time. For instance, LSTMs specifically address the challenge of long-term dependencies in sequences—an area where earlier neural models faltered. By effectively remembering information for long periods, LSTMs can detect complex attack patterns that unfold slowly, which are common in sophisticated network breaches. This capability makes them particularly valuable in scenarios where attackers employ low-and-slow tactics to evade detection. However, despite their strengths, they suffer from decreased performance as sequences lengthen. Given the inherently sequential nature of network traffic data, new sequential AI models, like the Transformer network model [11, 12], present a logical solution to this issue.
Transformers and their variants, leveraging self-attention mechanisms [11], have excelled in natural language processing (NLP) tasks, including text classification, dialogue recognition, and machine translation. Turning to the application of Transformer-based models such as convolution neural network (CNN) Transformer and vision Transformer (ViT) in network data analysis, these models bring the advantage of deep contextual understanding. Unlike RNNs and LSTMs, which process data linearly, Transformers analyze entire blocks of text in one go, capturing intricate patterns and dependencies across much longer sequences than was previously possible. This ability stems from their attention mechanisms, which weigh the importance of each word in the context of all others in the sequence. Inspired by the Transformer’s prowess in handling ordered data sequences, researchers have explored its applicability in intrusion and anomaly detection, demonstrating its effectiveness in various scenarios. Since network intrusion typically unfolds over time, most existing models lack the capability to capture time-series features, resulting in information loss. The attention mechanism of the Transformer, however, can effectively learn the temporal correlation of network intrusion data, thereby enhancing the accuracy of NIDS. The potential application of these advancements lies in their utilization with various cyber-data, including computer logs and network packets, which can be structured as text and conceptualized as distinct languages. A modest yet expanding body of literature has showcased successful applications of NLP to cyber data, such as identifying red team attacks within authentication logs. Once attacks are accurately identified, straightforward responses like quarantining or shutting down affected systems become feasible. Furthermore, leveraging windows event logs enables each computer to autonomously execute its response without necessitating feedback from other network sources. This autonomy is especially valuable in tactical networks, where connections may be unreliable and exhibit low throughput [13].
Transformers are a main component of large language models models due to their ability to capture complex dependencies in sequential data. For example, generative pre-trained transformer (GPT) [14] and bidirectional encoder representations from transformers (BERT) models [12] use Transformers to generate coherent and contextually relevant text. By leveraging attention mechanisms, Transformers enable LLM systems to understand and manipulate data at a more nuanced level, leading to more realistic and diverse outputs in tasks such as text generation, image generation, and speech synthesis [15]. The deployment of LLM in IDS leverages these contextual insights to a greater extent. By training on vast datasets, LLMs develop a nuanced understanding of what constitutes normal and anomalous network behavior. This DL capability enables them to identify even the most subtle signs of network intrusions, often before they fully manifest, thus providing a proactive rather than reactive security measure. In practical terms, this means that GPT and BERT can discern subtle anomalies in network traffic and log files by evaluating the data within the broader context of network behavior. For example, GPT’s generative capabilities allow it to predict potential future sequences based on past data, offering insights into expected versus actual network activity. Similarly, BERT’s bidirectional approach provides a holistic view of both past and future contexts within a dataset, enhancing its ability to pinpoint anomalies that would otherwise go unnoticed in a solely forward-feeding analysis. Thus, the LLM could be adapted at learning the benign patterns found within network traffic sequences, making it well-suited to detect alterations in network traffic patterns induced by even a minimal number of attacks [16].
@@
This survey aims to explore the cutting-edge IDS based on Transformers and LLM advances in AI techniques, which offer promising alternatives to address the weaknesses of existing ML and DL-based IDS state-of-the-art (SOTA) strategies. By reading this survey, the author aims to assist researchers in understanding the concept of enhancing IDS with Transformers and LLMs, as well as shedding light on future directions and perspectives.
@@ 1.1 Existing reviews and our contributions
  • The survey offers an extensive taxonomy of different Transformer models applicable to strengthen various IDS techniques, including Attention, CNN/LSTM-Transformer, ViT, GAN-Transformer, GPT, BERT and their variants, among others.
  • It presents a detailed classification of various Attention mechanisms and their possible configurations with traditional DL models, such as CNN.
@@ 1.3 Survey methodology
To identify and review existing studies on Transformers and LLM-based IDS, a comprehensive search was conducted across several leading publication databases renowned for their high-quality scientific research. The primary search was carried out in Scopus, which systematically includes databases such as Web of Science, Elsevier, IEEE, ACM Digital Library, Wiley, and IET Digital Library, among others.
Articles published between 2017 and 2024 were given priority. However, older publications were also considered when necessary to provide a historical context, dataset, metrics, etc. The survey focused on computer science and engineering studies from databases including IEEE Xplore, ScienceDirect, Wiley, Springer, and Taylor & Francis. Additionally, considering the recent and trending nature of the topic, high-quality pre-prints from arXiv, SSRN, and TechRxiv were selected. Only articles written in English were included in the final analysis. Key search terms used in the abstract, article title, and keywords. To efficiently manage the search and selection of relevant papers, the following query was structured to capture articles with focused relevance to Transformers and LLM-based IDS:
@@
The symbols & and || signify the logical operations AND and OR, respectively. The construction of this query is instrumental in refining the selection process, ensuring that only the most relevant studies are considered. This targeted approach has necessitated the development of explicit inclusion and exclusion criteria, detailed as follows:
Explicit inclusion criteria:
  • Research explicitly discussing Transformers or LLMs in the context of NLP, Cyberthreat, or IDS.
  • Studies must be published in indexed journals or conferences, ensuring they meet established academic standards and visibility.
Explicit exclusion criteria:
  • Studies that do not focus on the specified technological applications or topics.
  • Research ambiguously referencing "Transformers" in non-computational contexts, such as electrical engineering.
These inclusion and exclusion criteria are aimed at reinforcing the scientific rigor of the review process, ensuring that only pertinent and high-quality studies are considered. The final number of selected papers on IDS based purely on Transformers and LLMs is 118. Figure 1 provides more details about the involved research papers in terms of paper type, domain conducted, and distribution of papers based purely on Transformers and LLMs. It is obvious that Transformers and LLM-based IDS has emerged as a recent and rapidly growing field, with a notable surge in publications starting from 2022. Moreover, the inclusion of 40 pre-prints among the 180 cited papers underscores the trending nature and active research pursuit in this domain.
Figure 1: Bibliographic Statistics: (a) Analysis of research field and article type distribution from all references (180 papers). (b) Statistical distribution of 118 research papers on Transformers and LLMs for IDS (2017-2024).
@@ 2.2 Attack categories, datasets, and metrics
In-Vehicle 4 8.69 M 1:11 Flooding, fuzzy, Spoofing, Replay Yes AV IDS [37, 38] Yes [^2] https://ieee-dataport.org/open-access/car-hacking-attack-defense-challenge-2020-dataset
@@
Web server log 9 37,693 1:6 Suspicious and dangerous activities Yes Web IDS [57] Yes [^17] https://ieee-dataport.org/open-access/apache-web-server-access-log-pre-processing-web-intrusion-detection
@@ 3 Transformers-based IDS methods
Self-Attention Focuses on relevant parts of the input data for anomaly detection Improves detection by focusing on the most important parts of the data, increasing the efficacy of anomaly identification Can be computationally expensive and may require extensive tuning, posing challenges in dynamic network environments
Multi-head
Attention
Enhances self-Attention by using multiple Attention heads for better anomaly detection Provides multiple perspectives on the data, improving detection robustness, thus reducing false positives and negatives in IDS Increases model complexity and computational requirements, requiring significant resources for optimal performance
CANINE Attention Applied for deep analysis of of unsegmented sequential data packets in network traffic Effectively manages diverse, unsegmented data types, significantly enhancing robustness in anomaly detection. Requires extensive pre-processing of data, leading to increased operational overhead.
Local (hard)
Attention
Focuses on critical segments of data streams for quick anomaly detection in high-speed networks Enhances computational efficiency by focusing on key data points May miss anomalies outside the focused regions, limiting detection scope
Cross-Attention Employs in multi-modal IDS systems to analyze varied data sources like logs and network flows Enables thorough cross-referencing of data sources, improving detection accuracy Integration complexity and multi-source data management could be cumbersome
Conditional
Attention
Utilized in scenarios where attention is modulated based on external conditions in IDS Adapts dynamically to changing network environments, improving detection accuracy Complex integration with existing IDS frameworks may present challenges
CNN-Transformer Used for combining spatial feature extraction with sequential modeling Captures both local and global patterns in data, specifically enhancing anomaly detection in IDS Computationally intensive due to combining two complex models, potentially slowing real-time analysis
Vision Transformer Adapts vision-based Transformer models for network traffic visualization and anomaly detection Excellent at capturing intricate patterns in high-dimensional data, allowing for detailed anomaly visualization High computational cost and requires large datasets for training, which may be impractical for many organizations
GAN-Transformer Combines generative adversarial networks with Transformers for detecting and simulating anomalies Generates realistic anomalies, enhancing detection robustness, making the IDS adaptive to evolving threats High complexity and requires extensive training time, which may delay deployment and updates
BERT Applied for analyzing bidirectional network traffic data for anomaly detection Leverages forward and backward processing context for a better understanding of data patterns, enhancing precision in identifying network anomalies Requires large amounts of data for fine-tuning and significant computational resources, limiting its applicability in smaller-scale environments
RoBERTa Fine-tuned for anomaly detection in network traffic using robustly optimized BERT approach Improved performance over BERT due to enhanced training techniques, resulting in more accurate anomaly detection Even higher computational requirements and data needs compared to BERT, making it less suitable for resource-constrained settings
GPT Generates responses based on network behavior patterns for identifying anomalies Effective in generating realistic sequences, useful for simulation and detection, thereby aiding in proactive IDS measures May struggle with understanding context in bidirectional sense compared to BERT, leading to potential gaps in detection capabilities
3.1 Taxonomy of Attention Transformer
@@
In this mechanism, $Q$, $K$, and $V$ represent queries, keys, and values, respectively. Queries are used to direct the Attention mechanism’s focus. Keys are part of the input data used to calculate Attention weights, indicating the relevance of different parts of the input data. Values are the actual content that, after the application of Attention weights, are aggregated to form the output of the Attention mechanism. Additionally, $d_{k}$ denotes the dimension of the keys. This approach enables the model to concentrate on the most relevant parts of the data, thereby enhancing the precision and efficiency of security systems through more accurate threat identification and reduced false positives. Figure 4 summarizes the existing Attention categories, their roles in building various types of Transformers, and their related IDS applications.
Figure 4: Visual guide to attention types in seq2seq models, illustrating how focus shifts with input and output, along with their respective related works. The primary components of self-attention (used in [30, 58, 59, 60, 61, 62]) and its respective MHA (used in [6, 7, 40, 47, 63, 64, 65, 60, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]) are scaled dot-product and global (soft) attentions. Hierarchical attention constructs multiple layers of attention. Local hard attention is the main component of Sparse Transformer and Linformer. Cross-attention is a key component of encoder-decoder models (used in [83]), and ViT Transformers. Conditional attention (used in [84, 85]) is central to many LLM models such as BERT, GPT variants, and T5. CANINE attention is unique to the CANINE Transformer (used in [86]), where tokenization principles are not adhered to or realized.
@@ 3.3 CNN/LSTM-Transformer-based methods
A CNN-Transformer combines the strengths of CNNs and Transformers, two powerful architectures in DL. CNNs excel in processing structured grid data like images, efficiently capturing local dependencies through convolutional operations and detecting features at various levels of abstraction, making them pivotal in computer vision (CV) tasks. Transformers, on the other hand, leverage self-attention mechanisms to model long-range dependencies, excelling in handling sequential data such as text and time series. The integration of CNNs and Transformers enables more robust learning, particularly in tasks that require attention to both local and global data structures, such as object detection.
@@
The SOTA configurations of CNN-Transformers used in the realm of IDS are illustrated in Figure 6. Particularly, the scheme [86] alternates between CNN for spatial data processing and the Transformer for character-level text interpretation, following the configuration of Figure 6 (a), enhancing anomaly detection through deep contextual understanding and efficient feature extraction. The configuration depicted in Figure 6 (b) has been explored in numerous schemes. For instance, feature subsets covering multiple spaces are developed in [87] using CNN to enhance the spatial distribution of samples. Subsequently, the Transformer is employed to establish connections between features and to identify essential attributes, including the temporal and detailed aspects of the features, culminating in the successful detection of intrusion activities. Similarly, the work in [88] presents a hybrid neural network model aiming to improve feature extraction and detection effectiveness in traffic analysis. It combines dense CNNs and Transformers for feature fusion and time sequence extraction. Results on the CIC-IDS2018 dataset show 98% accuracy, surpassing existing models in performance metrics. Moving on, the study referenced in [89] employs 1D CNN and time series Transformer architectures to address the challenges posed by RNNs related to computational complexity and detection performance, particularly due to issues of information loss. The performance assessment on the CICDDoS2019 and CSE-CIC-IDS2018 datasets showed outstanding results, with the evaluation metrics mostly falling between 98.07% and 99.99%.
All the methods in [7, 90, 91, 92] are Transformer-based IDS approaches where the Transformer precedes the CNN block, as depicted in the configuration of Figure 6 (c). In the DDosTC model [7], the Transformer, followed by a sequence of CNN, work jointly to enhance DDoS attack detection. The Transformer processes sequential data, focusing on contextual dependencies, while the CNN extracts spatial features, optimizing for pattern recognition in network traffic. This synergy improves accuracy and robustness, outperforming models that rely solely on one type of network architecture. However, The DDosTC model may have limited generalizability due to testing on a single dataset and potential inefficiencies in real-time application scenarios with varied attack vectors. Similarly, In the IDS-INT model [90], Transformers first extract detailed feature representations from network interactions, focusing on semantic relationships within the data. After feature extraction, the synthetic minority oversampling technique (SMOTE) is applied to balance the dataset by synthesizing new samples for minority classes. Subsequently, the CNN processes these balanced features to deeply analyze and extract spatial patterns. This sequential processing significantly enhances the model’s ability to detect a wide range of network intrusions effectively. However, the method’s high computational demand could limit its applicability in resource-constrained environments. Yin et al. [91] proposed a mixed attention mechanism that enhances feature extraction by prioritizing influential attributes, while the CNN handles the spatial relationships between these features. This combination significantly boosts the model’s ability to discern subtle anomalies. Nevertheless, the capsule network’s dynamic routing mechanism is slower than traditional networks, requiring efficiency improvements.
@@
Figure 6: Various configurations of CNN-Transformer mechanisms include: (a) CNN and Transformer alternated [86]; (b) A sequence of CNN followed by a single Transformer [93, 87, 73, 54, 94, 95, 96]; (c) Transformer followed by sequence of CNN [7, 90, 91, 92]; (d) One-by-one CNN-Transformer feature fusion [43]; (e) A sequence of CNN followed by the last CNN and a single Transformer feature fusion [97, 67]; (f) Sequence of CNN and Transformer feature fusion [68].
The CNN and Transformer in the method proposed by Li et al. [43] follow the configuration shown in Figure 6 (d). This method enhances anomaly detection by capturing multi-scale features and optimizing sequential dependencies. Skip connections improve training stability and feature utilization, further boosting the model’s accuracy and robustness in time series data. The configuration shown in Figure 6 (e) has been adopted by [97, 67], where both employed CNN combined with MHA Transformer to enhance feature extraction by capturing both local and global dependencies in the data. Skip connections facilitate better gradient flow, preventing vanishing gradients and enabling deeper network training. This integration leads to more accurate and robust detection of anomalies or network intrusions in complex datasets. However, in [97], the hierarchical structure and deep architecture may result in increased computational complexity, potentially limiting real-time application capabilities, while in [67], the reliance on extensive network depth and MHA mechanisms can lead to overfitting, particularly with imbalanced or insufficiently diverse datasets. The configuration depicted in Figure 6 (f), where a sequence of CNN layers is fused with Transformer features, has already been adopted by Li et al. in [68]. The fusion of CNN and Transformer encoder captures both spatial and contextual features, significantly enhancing the precision and robustness of encrypted malicious traffic classification. However, the model’s complexity and resource-intensive pre-training increase computational demands, potentially limiting scalability and hindering deployment in real-time environments.
@@
Researchers have proposed numerous SOTA methods that combine CNN with LSTM to form hybrid schemes like CNN-LSTM, applied in various research fields including IDS [98], among others. These combinations have proven effective in enhancing overall accuracy. However, to handle long dependencies, researchers have increasingly either added Transformers or replaced LSTMs with Transformers. Wang et al. [40] suggest a model for detecting intrusions by integrating the pre-trained ResNet model, MHA-based Transformer, and bidirectional LSTM models into a single framework that captures both spatial and temporal characteristics of network traffic, leveraging their robust capabilities in learning data representations. This approach significantly reduces the time required to process complex, high-dimensional intrusion data through DL techniques. The objective is to address challenges in IDSs, such as inadequate feature extraction and imprecise classification with complex, nonlinear, and high-dimensional data. To ensure a balanced dataset, the SMOTE has been utilized for preprocessing.